# Entropy of self-avoiding branching polymers: mean field theory and Monte Carlo simulations

### v1

@@ Abstract
S4 Monte Carlo algorithm for spanning trees
As explained in main text, the Amoeba algorithm Seitz and Klein (1981); Rosa and Everaers (2016a, b); van der Hoek, Rosa, and Everaers (2024) is not suitable for simulating lattice trees when the volume fraction is $=1$. Therefore, in the particular case of spanning trees, we have devised a new Monte Carlo algorithm to sample trees based on the canonical distribution Eq. (10) in main text, therefore in the ensemble where nodes’ functionality is restricted to $\leq 3$ and such that traditional enumerating algorithms like Broder (1989); Aldous (1990) are not applicable.
@@
    Choose randomly a “$f=1$”-node.
@@
S5 Generalization to poor-solvent conditions
The SAT model considered in this work applies to polymers in so called good-solvent conditions Rubinstein and Colby (2003), namely to the situation where the effective interaction between the polymer and the surrounding solvent dominates over monomer-monomer (i.e. node-node) interactions. Following a procedure similar to the one presented in Sec. S1, it is however possible to obtain the mean field expression for the free energy of the system also in the opposite situation of poor-solvent conditions Rubinstein and Colby (2003), namely when an effective attractive interaction between nodes occupying nearest-neighbor lattice sites is considered in the system.
Indicating with $N_{mm}$ the total number of these node-node neighboring pairs and with $\epsilon>0$ the intensity of the attractive interaction, the energy associated with a given configuration $\mathcal{C}$ is written as $-\epsilon N_{mm}({\mathcal{C}})$. With reference to Eq. (S1), the corresponding grand canonical partition function of the system, denoted with $\Xi_{ps}$, is then given by
@@
where $\psi({\mathbf{x}})$ is a scalar field in correspondence of site $\mathbf{x}$ and ${\mathcal{D}}\psi=(2\pi)^{-V/2}(\det{\Delta})^{-1/2}\prod_{\mathbf{x}}\psi({\mathbf{x}})$ is the associated measure (notice that the introduction of a scalar field is a standard technique to count the number of nearest-neighbor interactions in lattice models, see Doniach, Garel, and Orland (1996); Marcato et al. (2023, 2024)). Starting from Eq. (S42), the procedure to obtain the mean field expression for the free energy of the system is exactly the same as the one presented in Sec. S1. Finally, the resulting free-energy per lattice site, which we denote with $\omega$, reads:
@@
Interestingly, there are two immediate and non-trivial consequences from Eq. (S43). First, it is clear that the branching probability $\langle\phi_{f}\rangle$ is independent of the interaction $\epsilon$, and then equal to the result in good-solvent presented in main text. Second, the phase diagram for the collapse transition of a self-avoiding branching polymer is predicted at the mean field level to be no different from the one of a linear chain, a feature that was already observed in previous works Nemirovsky, Dudowicz, and Freed (1992); De Los Rios, Lise, and Pelizzola (2001): this is easily seen as all linear terms (including the branching term) cancel by imposing the known condition for the collapse transition, $\partial^{2}\omega/\partial\rho^{2}=0$.
@@
Figure S3: SAT’s in $d=2,3,4$ spatial dimensions (panels from left to right). Mean fraction of branch-nodes, $\langle\phi_{3}\rangle$, as a function of the total number of nodes $N$. Dashed horizontal lines correspond to the mean field Eq. (9) in main text. Different colors/symbols are for different node fractions $\rho=0.0,0.1,0.2,0.4,0.8,1.0$ (same color/symbol codes as in Fig. 3 in main text), panels from top to bottom are for different values of the branch chemical potential $\beta\mu_{3}=-1,0,1,2$ (see legend). The filled square ($\blacksquare$) symbols for $d=3$ denote results for computer simulations of ideal lattice trees (i.e. no excluded volume).

### v2

@@ Abstract
S5 Monte Carlo algorithm for spanning trees
As explained in main text, the Amoeba algorithm Seitz and Klein (1981); Rosa and Everaers (2016a, b); van der Hoek, Rosa, and Everaers (2024) is not suitable for simulating lattice trees when the volume fraction is =1=1. Therefore, in the particular case of spanning trees, we have devised a new Monte Carlo algorithm to sample trees based on the canonical distribution Eq. (10) in main text, therefore in the ensemble where nodes’ functionality is restricted to $\leq 3$ and such that traditional enumerating algorithms like Broder (1989); Aldous (1990) are not applicable.
@@
    Choose randomly a “f=1f=1”-node.
@@
S6 Generalization to poor-solvent conditions
The SAT model considered in this work applies to polymers in so called good-solvent conditions Rubinstein and Colby (2003), namely to the situation where the effective interaction between the polymer and the surrounding solvent dominates over monomer-monomer (i.e. node-node) interactions. Following a procedure similar to the one presented in Sec. S2, it is however possible to obtain the mean-field expression for the free energy of the system also in the opposite situation of poor-solvent conditions Rubinstein and Colby (2003), namely when an effective attractive interaction between nodes occupying nearest-neighbor lattice sites is considered in the system.
Indicating with $N_{mm}$ the total number of these node-node neighboring pairs and with $\epsilon>0$ the intensity of the attractive interaction, the energy associated with a given configuration $\mathcal{C}$ is written as $\epsilon N_{mm}({\mathcal{C}})$. With reference to Eq. (S2), the corresponding grand canonical partition function of the system, denoted with $\Xi_{ps}$, is then given by
@@
where $\psi({\mathbf{x}})$ is a scalar field in correspondence of site $\mathbf{x}$ and ${\mathcal{D}}\psi=(2\pi)^{-V/2}(\det{\Delta})^{-1/2}\prod_{\mathbf{x}}\psi({\mathbf{x}})$ is the associated measure (notice that the introduction of a scalar field is a standard technique to count the number of nearest-neighbor interactions in lattice models, see Doniach, Garel, and Orland (1996); Marcato et al. (2023, 2024)). Starting from Eq. (S49), the procedure to obtain the mean-field expression for the free energy of the system is exactly the same as the one presented in Sec. S2. Finally, the resulting free-energy per lattice site, which we denote with $\omega$, reads:
@@
Interestingly, there are two immediate and non-trivial consequences from Eq. (S50). First, it is clear that the branching probability $\langle\phi_{f}\rangle$ is independent of the interaction $\epsilon$, and then equal to the result in good-solvent presented in main text. Second, the phase diagram for the collapse transition of a self-avoiding branching polymer is predicted at the mean-field level to be no different from the one of a linear chain, a feature that was already observed in previous works Nemirovsky, Dudowicz, and Freed (1992); De Los Rios, Lise, and Pelizzola (2001): this is easily seen as all linear terms (including the branching term) cancel by imposing the known condition for the collapse transition, $\partial^{2}\omega/\partial\rho^{2}=0$.
@@
Figure S4: SAT’s in $d=2,3,4$ spatial dimensions (panels from left to right). Mean fraction of branch-nodes, $\langle\phi_{3}\rangle$, as a function of the total number of nodes NN. Dashed horizontal lines correspond to the mean-field Eq. (9) in main text. Different colors/symbols are for different node fractions $\rho=0.0,0.1,0.2,0.4,0.8,1.0$ (same color/symbol codes as in Fig. 3 in main text), panels from top to bottom are for different values of the branch chemical potential $\beta\mu_{3}=-1,0,1,2$ (see legend). The filled square ($\blacksquare$) symbols for $d=3$ denote results for computer simulations of ideal lattice trees (i.e. no excluded volume).